OpenStudy (anonymous):

which expression is equivalent to 2(m - 5) + 6 2m - 1 2m + 22 2m - 4 2m + 16

Nnesha (nnesha):

distribute parentheses by outside number here is an example \[\huge\rm \color{red}{a}(b+c)= \color{Red}{a} \times b \color{blue}{+} \color{Red}{a} \times c = \color{Red}{a}b\color{blue}{+}\color{Red}{a}c\]